OpenStudy (anonymous):

Give lengths to three significant digits and angle measures to the nearest tenth of a degree or nearest ten minutes. Solve each right triangle ABC.

OpenStudy (anonymous):

angle B= 65.4 degrees, a=2.35

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

I see. What's angle A, then?

OpenStudy (anonymous):

angle A is 24.6 degrees

terenzreignz (terenzreignz):

mhmm. so... we need to find the length of side b. How do we do that?

OpenStudy (anonymous):

i don't know is it ...\[\sin 65.4=\frac{ 2.35 }{ c }\]

terenzreignz (terenzreignz):

In fact, that equation is true. Go ahead, then, find c.

OpenStudy (anonymous):

c=2.6?

terenzreignz (terenzreignz):

good enough. now find b.

OpenStudy (anonymous):

b is 1.1?

terenzreignz (terenzreignz):

how?

OpenStudy (anonymous):

\[\cos65.4=\frac{ b }{ 2.58 }\]

terenzreignz (terenzreignz):

I'd rather you didn't use the hypotenuse. Because you already did some rounding here. Try to exclusively use the given info as much as you can. Take advantage of the tan function ^_^

OpenStudy (anonymous):

i keep doing that lol... bad habit

terenzreignz (terenzreignz):

tan A = a/b And you already angle A, side a, so find b :)

OpenStudy (anonymous):

i get the same thing.. :/

terenzreignz (terenzreignz):

You should, because what you did earlier was correct, in essence. I'd just rather you arrive at the answer using the given info... I mean, heaven forbid, what if you got a wrong answer for c, huh? ;)

OpenStudy (anonymous):

My answer key says that c=5.65 and b=5.14... so I don't why..

terenzreignz (terenzreignz):

Hang on.

terenzreignz (terenzreignz):

Okay, clean slate, okay? Just for kicks, could you draw the right triangle?

OpenStudy (anonymous):